OpenStudy (anonymous):

what is the domain and range of f(x) = -3sin(x)?

OpenStudy (anonymous):

domain: [0,pie/2] range : [-3,+3]

OpenStudy (anonymous):

@dhashni, the domain of sine and cosine is all real numbers. there is no restriction

OpenStudy (across):

As satellite said, \(\sin(x)\) is defined everywhere.\[\]

OpenStudy (anonymous):

the graph of f(x) = sin(x) – 5?

OpenStudy (anonymous):

you have the range correct, however. since \[-1\leq \sin(x)\leq 1\] then \[-3\leq -3\sin(x)\leq 3\]

OpenStudy (anonymous):

then when we restrict its domain?

OpenStudy (anonymous):

\[y=\sin(x)-5\]looks exactly like the graph of \[y=\sin(x)\] only shifted down 5 units

OpenStudy (anonymous):

@rosy, it is your function, you can do what you like with it. if you choose to restrict the domain, you may. but as a convention we take the domain to be the largest possible set for which we get a real number output, and for sine and cosine that set is all real numbers

OpenStudy (anonymous):

i think when we want to find the inverses of sine and cosine functions then we restrict the domain from all rel nos. to [0,pie] isn't it?